Consider a set of three shares in different companies. On any day the market price of each share can either increase (I), decrease (D) or remains unchanged (U).Assuming P(D) =P(I) =P(D)
(i) Write the sample space for the behaviour of the 3 shares
(ii) Compute the probabilities of the following events.
A = ( shares 1 , 2 , 3 behaves in the same way )
B = ( shares 1 and 2 increase in value )
C = ( none of the share price increases )
1
Expert's answer
2020-08-06T17:41:41-0400

"Increase =I"

"Decrease=D"

"Unchanged=U"

solution i)


sample spaces:


"Share 1=\{I,D,U\}"

"Share2=\{I,D,U\}"

"Share3=\{I,D,U\}"


solution ii)


Let X be the movement from share1, Y be the movement from share2 and Z from share3


Therefore:



"P(X=x)=P(D)orP(U)or P(I)"


Since:



"P(D)=P(I)=P(U)"


"P(X=x)=\frac{1}{3}"

Similarly:


"P(Y=y)=P(Z=z)=\frac{1}{3}"


P(A) where:



"A=\{\{x\in share1\}=\{y\in share2\}=\{z \in share3\}\}"


"=(\frac{1}{3}×\frac{1}{3}×\frac{1}{3})×3=\frac{1}{9}"

answer: "\frac{1}{9}"


P(B) where:



"B=\{\{X=I\} and \{Y=I\} and\{Z=D or U\}\}"


"=\frac{1}{3}×\frac{1}{3}×(\frac{1}{3}+\frac{1}{3})=\frac{2}{27}"



answer: "\frac{2}{27}"


P(C) Where:



"C=\{\{X=DorU\}and\{Y=DorU\}and\{Z=DorU\}"

"=(\frac{1}{3}+\frac{1}{3})×(\frac{1}{3}+\frac{1}{3})×(\frac{1}{3}+\frac{1}{3})=\frac{8}{27}"

answer: "\frac{8}{27}"